# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### v2

@@ Abstract
To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST). LIST is featured with two novel components. Firstly, we propose a lightweight and effective relevance model that is capable of learning both spatial and textual relevance. Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index to accelerate the top-k search process of the proposed relevance model, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. There exist two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
@@ 1 Introduction
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to search for a ‘delicious Italian restaurant’ that is close to the user’s hotel.
@@
Spatial keyword queries have applications in many real-world scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions. To meet diverse user needs, various spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 . Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance. Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (formulated in Section 3.1). According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient . Most of the existing studies congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 on spatial keyword queries focus on improving the efficiency of handling spatial keyword queries. As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
@@
Although many indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited. As discussed earlier, TkQ typically uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue, e.g., synonyms that consist of different tokens may convey the same meanings, which limits their effectiveness (detailed in Section 3.2). For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models. Another limitation arises from using the linear function of distance to calculate spatial relevance. This practice is based on an assumption that the user’s spatial preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and does not hold on our real-life datasets, where the spatial preference for geo-textual objects exhibits a significant (non-linear) decrease with the increase of distance. This mismatch with real-world situations results in limited effectiveness (as to be detailed in Section 3.2).
@@
To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue. To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models. For example, the recently proposed DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings, identifies top-k relevant terms from object description for each query keyword based on the word embeddings, and utilizes an attention mechanism to aggregate the relevance scores of every query keyword as the textual relevance. It also uses the linear function of distance as the spatial relevance. The interactions between word embeddings bring very high query latency. And it relies on interactions between the word embeddings of queries and objects, which cannot be pre-computed until the query-object pairs are given, therefore there is no efficient index designed for it (detailed in Section 3.2).
To address the aforementioned limitations, this work aims to (1) design a lightweight yet effective relevance model, and (2) design an index to expedite the top-k search process of the proposed relevance model.
For the first objective, we propose a lightweight yet effective relevance model for embedding based spatial keyword queries, which capture the semantic meaning of query keywords in one high-dimensional embedding. The relevance model comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module. The dual-encoder module converts query keywords and object descriptions into two separate embeddings and evaluates textual relevance based on the two embeddings. To overcome the limitation of linear distance function, as explained earlier, we propose a new learning-based spatial relevance module, which is capable of learning from real-world datasets to better evaluate spatial relevance. Additionally, a weight learning module is proposed to adaptively learn a weight to balance the textual and spatial relevance.
@@
For the second objective, existing geo-textual indexes are developed based on traditional models, and cannot be used to handle embedding based spatial keyword queries. Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for embedding based spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section 5.2). An intuitive idea is to adapt these ANNS indexes to incorporate the spatial factor. For instance, the IVF index jegouProductQuantizationNearest2011 clusters embeddings using the K-means algorithm and route queries to a small number of close clusters to reduce the search space. However, incorporating spatial factors into the K-means algorithm is challenging as it is hard to set the weight to balance spatial and embedding similarities during index construction. Manually setting the weights is not only laborious but also ineffective, resulting in inferior effectiveness (as to be shown in Section 5.2). It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
@@
To fill this gap, we develop a novel machine learning based ANNS index. To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones. Existing learning-to-cluster studies Hsu16_KCL ; Hsu18_L2C demonstrate that high-quality pairwise similar/dis-similar labels are essential for training. Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse , and high-quality negative labels are absent. Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built on the cluster results. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
@@
Based on the two proposed techniques, namely the lightweight yet effective relevance model and the machine learning based ANNS index, we propose a new solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries (called LIST). LIST employs the proposed relevance model to pre-compute embeddings for geo-textual objects and uses the proposed index to cluster the relevant objects together while separating irrelevant ones. Given embedding based spatial keyword queries, LIST routes queries to a subset of clusters to reduce the search space. Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (detailed in Section 4).
@@
  • New Relevance Model. We develop a lightweight yet effective relevance model. It is featured with a dual-encoder module to convert query keywords and object descriptions into two separate embeddings and compute textual relevance based on the two embeddings, a novel learning-based spatial relevance module to learn from real-word datasets to better evaluate the spatial relevance, and a new weight learning mechanism to determine the weights between textual and spatial relevance.
@@
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets. Experimental results show that LIST significantly outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
@@ 2 Related Work
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient are proposed. Among them, Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance. However, these traditional methods have limited effectiveness.
@@
Different from previous spatio-textual relevance models that rely on word embeddings and complex models to compute textual relevance, our proposed model employs a dual encoder module to transform query keywords and object descriptions into two embeddings, computes textual relevance based on two embeddings, and we also propose a new learning-based spatial module to learn from real-world datasets to better estimate spatial relevance.
@@
[…]
@@
Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model, which relies on embeddings to compute textual relevance. It is still an open problem for geo-textual data.
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models guo2020deep . The first category of models (e.g., ARC-II DBLP:conf/nips/HuLLC14 and MatchPyramid DBLP:conf/aaai/PangLGXWC16 ) calculates the word-level similarity between queries and documents for textual relevance. DrW liu2023effectiveness belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM huang2013learning ) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of the dual-encoder architecture DBLP:conf/emnlp/KarpukhinOMLWEC20 ; DBLP:conf/icml/GuuLTPC20 , which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 . For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep . Note that, models in the first category like DrW liu2023effectiveness are not efficient for online computation and retrieval since the complex word interaction function cannot be pre-calculated until we see the input query-object pairs according to guo2020deep .

### v3

@@ Abstract
To tackle these issues, we consider embedding based spatial keyword queries, which capture the semantic meaning of query keywords and object descriptions in two separate embeddings to evaluate textual relevance. Although various models can be used to generate these embeddings, no indexes have been specifically designed for such queries. To fill this gap, we propose LIST, a novel machine learning based Approximate Nearest Neighbor Search index that Learns to Index the Spatio-Textual data. LIST utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. There are two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Additionally, we introduce a learning based spatial relevance model that can integrates with various text relevance models to form a lightweight yet effective relevance for reranking objects retrieved by LIST. Experimental results show that (1) our lightweight embedding based relevance model significantly outperforms state-of-the-art relevance models; (2) LIST outperforms state-of-the-art indexes, providing a better trade-off between effectiveness and efficiency.
@@ 1 Introduction
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data that contains both text and geographical location, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to search for a ‘delicious pizza restaurant’ that is close to the user’s location. A returned object could be a nearby restaurant named ‘Pizza Palace’.
@@
Spatial keyword queries have applications in many real-world scenarios such as geographic search engines DBLP:conf/sigmod/ChenSM06 , location-based services DBLP:conf/cikm/ZhouXWGM05 , and local web advertising tailored to specific regions dong2021continuous . To meet diverse user needs, various types of spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 . Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves the top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance. Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (as to be formulated in Section 3.1). According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient . Most of the existing studies congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 on spatial keyword queries focus on improving the efficiency of handling spatial keyword queries. As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
@@
Motivations. Despite various indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited. As discussed earlier, TkQ uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue DBLP:conf/emnlp/KarpukhinOMLWEC20 ; renRocketQAv2JointTraining2021 ; DBLP:conf/naacl/QuDLLRZDWW21 , e.g., synonyms that consist of different tokens may convey the same or similar meanings, which limits their effectiveness (as to be detailed in Section 3.2). For example, consider searching for an ‘Italian restaurant’ on a location-based app like Foursquare. Even if a nearby restaurant named ‘Pasta House’ exists, traditional models will not be able to retrieve this result because they cannot semantically match ‘Italian restaurant’ with ‘Pasta House’, it will receive zero textual relevance score.
@@
To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue. For instance, DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings and identifies top-k relevant terms from object description for each query keyword based on the word embeddings. It then uses an attention mechanism to aggregate the relevance scores between each keyword and its top-k relevant terms, determining the overall textual relevance. In our experiment, DrW takes over 7 seconds to answer a query in the Geo-Glue dataset using brute-force search. Moreover, to the best of our knowledge, there are no efficient indexes designed to expedite the top-k search process for these deep relevance models.
Objective and Challenges. To this end, we aim to develop an efficient index for deep learning models to answer spatial keyword queries. However, these models, such as DrW, often suffer from high latency and it is difficult to design indexes for them due to their complex word interaction functions, which rely on word embeddings to calculate the word-level similarity between query keywords and object descriptions for evaluating textual relevance. According to guo2020deep , these functions not only result in high latency but also cannot be pre-calculated until the query-object text pairs are seen, making it challenging to build an index for them.
To address this, we consider the embedding based spatial keyword queries, which capture the semantic meanings of query keywords and object descriptions into two separate embeddings and evaluate textual relevance based on the two embeddings. Many text models can be used for the embedding based spatial keyword queries, ranging from earlier Word2Vec mikolov2013efficient to the recent advancements in pre-trained language models zhao2024dense .
@@
The challenge lies in developing an index for embedding based spatial keyword queries. Existing geo-textual indexes are developed based on traditional models such as BM25 and TF-IDF, which cannot be used to handle embedding based spatial keyword queries. Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for embedding based spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section 5.2). An intuitive idea is to adapt these ANNS indexes to incorporate the spatial factor. For instance, the IVF index jegouProductQuantizationNearest2011 clusters embeddings using the K-means algorithm and route queries to a small number of close clusters to reduce the search space. However, incorporating spatial factors into the K-means algorithm is challenging as it is hard to set the weight to balance spatial and embedding similarities during index construction. Manually setting the weights is not only laborious but also ineffective, resulting in inferior effectiveness (as to be shown in Section 5.2). It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
@@
To fill this gap, we develop a new machine learning based ANNS index that is applicable to any relevance models for embedding based spatial keyword queries only if query and objects are represented by two embeddings. To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones. Existing learning-to-cluster studies Hsu16_KCL ; Hsu18_L2C demonstrate that high-quality pairwise similar/dis-similar labels are essential for training. Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse, and high-quality negative labels are absent. Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built based on the cluster results. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
@@
We call the proposed index as LIST, which learns to index spatio-textual data for answering embedding based spatial keyword queries. LIST employs a relevance model to pre-compute embeddings for geo-textual objects, and learns to index these embeddings. Given an embedding-based spatial keyword query, LIST routes the query to a subset of clusters to reduce the search space. Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (as to be detailed in Section 4).
In addition, we improve the spatial relevance for embedding based spatial keyword queries. Existing relevance models such as DrW typically use the linear function of distance, assuming that spatial preference decreases linearly with distance. However, this assumption was not examined previously. We find it does not hold on our real-life datasets, and that the spatial preference exhibits a significant (non-linear) decrease with the increase of distance (as to be detailed in Section 3.2). This motivates us to design a new spatial relevance model to fit this real-world pattern better. A straightforward approach would be to employ an exponential function of distance. However, our experiments indicate that this simple method is even less effective compared to the linear function (as to be shown in Section 5.6). Instead, we design a learning based spatial relevance model, which learns from real-world datasets to better evaluate spatial relevance. The spatial model is built on two real-world patterns: (1) spatial relevance increases as distance decreases, and (2) spatial relevance increases in a stepwise manner rather than continuously as distance decreases. Therefore, we designed the model as a learnable monotonic step function. Additionally, we introduce a weight learning module to adaptively learn a weight for embedding based spatial keyword queries to balance the textual and spatial relevance.
@@
  • New Spatial Relevance Model. We develop a novel learning based spatial relevance module, which is capable of learning from real-world datasets to better evaluate the spatial relevance. This model can integrate with various textual relevance models for embedding based spatial keyword queries.
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our solution LIST on three real-world datasets. Experimental results show that (1) our lightweight embedding based relevance model significantly outperforms the state-of-the-art relevance models for effectiveness by an improvement up to 31.60% and 59.92% in terms of NDCG@1 and Recall@10, respectively; (2) LIST outperforms existing state-of-the-art indexes on the three datasets, providing a better trade-off between effectiveness and efficiency.
@@ 2 Related Work
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient have been proposed. Among them, Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance. However, these traditional methods have limited effectiveness.
@@
Different from previous deep learning based relevance models that rely on word embeddings and complex models to compute textual relevance, we consider embedding based spatial keyword queries, which transform query keywords and object descriptions into two separate embeddings and computes textual relevance based on the two embeddings. We also propose a new learning based spatial module to learn from real-world dataset to better estimate spatial relevance.
@@
Several indexes DBLP:journals/www/QianXZZZ18 ; DBLP:journals/geoinformatica/ChenXZZLFZ20 have been introduced to incorporate semantic representations into the TkQ scheme. For example, S2R-Tree DBLP:journals/geoinformatica/ChenXZZLFZ20 projects the word embeddings to an $m$-dimensional vector using a pivot-based technique ($m$ as low as 2). Consequently, it employs R-trees to index objects based on their geo-locations and $m$-dimensional vectors hierarchically. When the dimensionality reaches hundreds, which are common for embeddings, such methods are no better than a linear scan due to the curse of dimensionality indyk1998approximate .
@@
Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model that is developed for embedding based spatial keyword queries. It is still an open problem for geo-textual data.
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models guo2020deep . The first category of models (e.g., ARC-II DBLP:conf/nips/HuLLC14 and MatchPyramid DBLP:conf/aaai/PangLGXWC16 ) calculates the word-level similarity between queries and documents for textual relevance. DrW liu2023effectiveness belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM huang2013learning ) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of Pre-trained Large Language Models (PLMs), which extract global semantic representation from textual content, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 . For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep . Note that, models in the first category like DrW liu2023effectiveness are not efficient for online computation and retrieval since the complex word interaction function cannot be pre-calculated until we see the input query-object pairs according to guo2020deep .